OpenStudy (anonymous):

Twice last month, Judy rented a car from a car rental company and traveled around the southwest on business. The company rents its car for a daily fee plus an additional charge per mile driven. Judy recalls that her first trip lasted 4 days, she drove 450 miles and the rental car costed her $222. On her second business trip she drove 200 miles in 3 days and paid $139 for the rental. Find the daily fee and find the mileage charge.

OpenStudy (anonymous):

We can set up a system of equations. let d=daily fee and let m=cost per mile use the elimination method. 4d+450m=222 <------*-3 3d+200m=139 <---------*4 ---------------- -12d-1350m=-666 12d+800m=556 add them together, d cancels out. -550m=-110 m=.2 Substitute m into original equation and solve for d. 4d+450(.2)=222 4d+90=222 4d=132 d=33 So daily charge is $33 and cost per mile is $.20